A student has taken two math tests. On the first, he received a score of 46 out of 50 questions correct, and on the second, he received a score of 27 out of 30 questions correct. What is the experimental probability that the student answered a question correctly?

Answer by ewatrrr(24785)

1st: p(correct) = 46/50
2nd: p(correct) = 27/30
both: p(correct) = 73/80
P(a particular question was correct) = 73/80